OpenStudy (anonymous):

Explain how the trinomial 2x2+ 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping

OpenStudy (lukecrayonz):

(x+5)(2x+3)

OpenStudy (anonymous):

(2x+3)(x+5)

OpenStudy (anonymous):

rewrite the middle term, 13x, as the product of the lead coefficient and the constant, which is 30; thus 3*10=30 and 3+10=13\[2x^2+10x+3x+15\]then you can factor by grouping\[=2x(x+5)+3(x+5)\]\[=(x+5)(2x+3)\]